# On an Optimal Stopping Problem with a Discontinuous Reward

### v2

@@ Abstract.
Keywords: Variable annuities, surrender option, optimal stopping, American options, optimal surrender boundary, early exercise premium, free-boundary value problem, variational inequalities, continuous-time Markov chain approximation.
JEL Classifications: C63, G12, G13, G22
${}^{\ast}$orresponding author. Email: Anne.MacKay@USherbrooke.ca
@@ 1. Introduction
Variable annuities (VA) are structured products sold by insurance companies that are mainly used for retirement planning. An initial premium is deposited in an investment account (or fund) whose return is linked that of one or more risky assets. They are similar to mutual funds, but they also offer financial guarantees at the end of a pre-determined accumulation period. The embedded guarantees are akin to long-dated options, but they are funded by a periodic fee, generally set as a percentage of the investment account value, rather than being paid upfront. In this paper, we focus on a guaranteed minimum maturity benefit (GMMB), which provides the policyholder a minimum guaranteed amount at maturity of the contract. Another common feature of variable annuities contracts is that policyholders generally have the right to surrender, or lapse, their contract prior to maturity. When policyholders choose to do so, they receive the value accumulated in the investment account, reduced by a penalty for early surrender (or surrender charge). The uncertainty faced by VA providers with respect to early termination is known as surrender risk. Early surrenders can entail negative consequences such as liquidity issues. Thus, incorporating adequate surrender assumptions in the pricing of VAs is essential for risk management purposes; see Niittuinperä (2022).
@@
Different early surrender modelling approaches have been proposed in the literature (see Bauer et al. (2017) and Feng et al. (2022) for a review), from the use of utility functions, Gao and Ulm (2012), to modern statistical techniques, Zhu and Welsch (2015). Early surrenders are also often expressed as a decision taken by policyholders on a strictly rational basis, meaning that the contract is terminated as soon as it is optimal to do so from a financial perspective (see Bernard et al. (2014a), Jeon and Kwak (2018), Kang and Ziveyi (2018), MacKay et al. (2023), and Milevsky and Salisbury (2001), among others). Bauer et al. (2017) discusses the impact of other factors on policyholder behavior, such as taxes and expenses. This article gave rise to another stream of literature in which market frictions such as taxation rules are considered (Alonso-García et al. (2022), Bauer and Moenig (2023), and Moenig and Bauer (2016)), which helps to explain the discrepancies between market and model fee rates, Moenig and Bauer (2016). Other recent studies include lapse and reentry strategies in their analysis (Bernard and Moenig (2019) and Moenig and Zhu (2018)), whereas Moenig and Zhu (2021) consider a third-party investor to whom the policyholder can sell her contract.
@@
The goal of this paper is to study the properties of the value function of the VA contract under the assumption that the policyholder maximizes its risk-neutral value, when the fee and the surrender penalty (or charge) are both time and state-dependent (that is, depending on the value of the VA account, see Bernard et al. (2014a)). Our setup includes, among others, the constant fee case, the state-dependent fees of Bernard et al. (2014a), Delong (2014), and MacKay et al. (2017), and the time-dependent fees of Bernard and Moenig (2019) and Kirkby and Aguilar (2023). Under this assumption, valuing a variable annuity contract is equivalent to solving an optimal stopping problem similar to pricing an American option. However, the financial guarantee being applied only at maturity in a VA contract creates a discontinuity in the reward function at maturity, which sets it apart from the continuous reward function of the American option and complicates the optimal stopping problem involved in the pricing of VAs. The impact of the surrender charge and the fee structure on the optimal surrender strategy has been studied in the Black-Scholes framework by Bernard et al. (2014a), Bernard and MacKay (2015), Bernard and Moenig (2019), MacKay (2014), MacKay et al. (2017), and Moenig and Zhu (2018), and in a more general setting by Kang and Ziveyi (2018) and MacKay et al. (2023). However, these articles approach the problem from a numerical perspective. To the author’s knowledge, it is the first time that such an extensive analytical study of the optimal stopping problem involved in the pricing of a variable annuity contract is performed in the literature.
@@
Recently, Luo and Xing (2021) studied variable annuity contracts under regime-switching volatility models from an optimal stopping perspective. However, they consider a different surrender benefit, which results a continuous reward function similar to that of a put option. Chiarolla et al. (2022) perform an analysis similar to ours, but for participating policies with minimum rate guarantee and surrender option. Participating policies are akin to variable annuities in that the premium paid by the policyholder tracks a financial portfolio, subject to a minimum rate guarantee. However, the problem they consider is fundamentally different from ours in multiple ways. i) In variable annuity contracts, guarantees are funded via periodic fees set as a percentage of the sub-account value. This creates a discrepancy between the fee amount and the value of the financial guarantee, which becomes an incentive for the policyholder to surrender the policy early (Milevsky and Salisbury (2001)). In participating policies, there is no such ongoing fee; upon termination or at maturity, the policyholder receives the reserve and a given percentage of the surplus, defined as a percentage of the tracking portfolio over the policy reserve (the intrinsic value). ii) In variable annuities, the surrender value differs from that of the maturity payout, which creates a time discontinuity in the reward function at maturity; whereas in participating policies, the intrinsic value is paid at maturity or upon early termination, so that the reward function is independent of time and continuous. iii) In participating policies, the contract is terminated if the underlying portfolio falls below the reserve. There is no such feature in a variable annuity, that is, early termination is at the sole discretion of the policyholder. The resulting optimal stopping problem we study in this paper is thus very different from the one considered by Chiarolla et al. (2022), even if they are motivated by similar insurance products.
@@
The shape and the (non-)emptiness of the exercise region of an American option with a time-homogeneous reward function has first been studied rigorously by Villeneuve (1999). Other authors, such as Jönsson et al. (2006), Kotlow (1973), and Villeneuve (2007), also studied the conditions under which the exercise region has a particular shape. These results are however not directly applicable to the problem we study in this paper, because of the time-dependence and the discontinuity of our reward function. Under the assumption of a general fee and surrender charge structure, we are nonetheless able to identify a condition, expressed as a partial differential inequality, under which the optimal stopping time is always at maturity of the contract, thus mitigating surrender risk. This condition provides a powerful tool to understand the interplay between the fees and surrender charges, and its effect on the optimal stopping strategy. It can also be used to further study and characterize the shape of the optimal exercise, or surrender, region. In particular, we show that the surrender region can be a disconnected set and that the optimal surrender boundary can be discontinuous, thus illustrating a major difference between the pricing of VA contracts and standard American options. Our results extend the work of Milevsky and Salisbury (2001), who consider an infinite horizon, to a more general finite horizon problem.
@@
The main contributions of this paper are listed below.
@@
  • We present an alternative representation for the value function of the VA contract in terms of a continuous reward function. This representation is used to perform a rigorous theoretical study of the value function. Our results justify the use of various numerical methods already applied to VA pricing in the literature.
  • We provide two integral representations for the value function, which are used to further analyze the value of a VA contract and to characterize the (non-) emptiness of the surrender region. The continuation premium decomposition is new to the actuarial and American option pricing literature and allows for the development of new valuation methods for variable annuity contracts.
  • We obtain conditions under which the original optimal stopping problem and the one with an alternative continuous reward function are equivalent.
@@ 2.1. Market Model
On a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, let $W=\{W_{t}\}_{t\geq 0}$ be a standard Brownian motion whose augmented filtration is denoted by $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$. $\mathbb{Q}$ is the probability measure used to price assets presented below.
@@
where $r$, $\sigma$ are deterministic constants with $\sigma>0$, and $W=\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion. It is easy to verify that (1) has a unique strong solution given by $B_{t}=B_{0}e^{rt}$ and
@@
Most readers will recognize the so-called Black-Scholes model, presented here directly in terms of its unique risk-neutral measure.
@@ 2.2. Variable Annuity Contract
In this section, we describe a simplified variable annuity contract offering a guaranteed minimum accumulation benefit at maturity $T\in\mathbb{R}_{+}$. Throughout this paper, $\mathbb{R}_{+}$ denotes the strictly positive real numbers, that is +=$\mathbb{R}_{+}=\{x\in\mathbb{R}|x>0\}$. At inception of the contract, the policyholder deposits an initial premium in an investment sub-account (or fund) tracking the financial market. Here, we assume that this sub-account tracks the risky asset with price process $S$. We denote by $F=\{F_{t}\}_{0\leq t\leq T}$ the value process of the sub-account.
We remark that variable annuity policyholders usually benefit from a return-of-premium guarantee if they die before maturity. Under the assumption that mortality risk is completely diversifiable, it is straight-forward to add the death benefit in the analysis of the contract, see for example MacKay et al. (2017). For this reason, in this paper, we ignore mortality risk.
The guarantees embedded in VA contracts are funded via a fee levied continuously from the investment sub-account at a rate $c_{t}$ defined as
@@
where $C:[0,T]\times\mathbb{R}_{+}\rightarrow[0,1]$, so that (4) has a unique strong solution.This fee structure is general enough to include state-dependent (Bernard et al. (2014b), Delong (2014), MacKay et al. (2017)) and time-dependent fees (Bernard and Moenig (2019), Kirkby and Aguilar (2023)). The sub-account value is then defined by
@@
At maturity, the policyholder receives the maximum between a pre-determined amount $G\in\mathbb{R}_{+}$ and the value of the investment sub-account. Given $F_{t}=x$, the time-$t$ risk-neutral value of the maturity benefit is given by
@@
Should she decide to surrender her contract prior to maturity, the policyholder receives the amount accumulated in the sub-account subject to a penalty charge. We assume that this charge, expressed as a percentage of the sub-account, can depend on time and on the sub-account value. If no surrenders occur, the maturity benefit is paid at $T$.
@@
where $g:[0,\,T]\times\mathbb{R}_{+}\rightarrow(0,1]$ is in $C^{1,2}$, non-decreasing in $t$ and satisfies $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$. This function represents the amount received by the policyholder upon surrender at $t<T$ or at maturity $T$, given that the sub-account has value $x$.
In practice, $1-g(\cdot,\cdot)$ is often called the surrender charge. It is non-increasing in time; a later surrender will yield a higher proportion of the account value. In this work, we often use the term surrender charge function to refer to either $g(\cdot,\cdot)$ or $1-g(\cdot,\cdot)$.
Imposing $g(T,\cdot)=1$ is common in the actuarial literature. It allows the function $g(\cdot,\cdot)$ to be defined on the closed interval $[0,T]$ and represents the fact that at maturity, the policyholder is entitled to the full amount accumulated in the account. In practice, it is also common to see the surrender charge vanish before maturity. Examples of surrender charges are listed in Palmer (2006).
@@
When $x<G$, the function $t\mapsto\varphi(t,x)$ is discontinuous at $T$ since
@@
Under the assumption that the policyholder maximizes the risk-neutral value of her contract, its time-$t$ value is given by
(7) $v(t,\,x)=\sup_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}[e^{-r(\tau-t)}\varphi(\tau,F_{\tau}^{t,\,x})],$
where $\mathcal{T}_{t,\,T}$ is the set of all stopping times taking value in the interval $[t,\,T]$.
The difference between the value of the full variable annuity contract and the present value of the maturity benefit is the value of the surrender right, which we denote by $e:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and define as
@@ 3. Optimal Stopping Time
In this case, we also say that $\tau^{\star}$ is optimal for (8). For a general reward function $\varphi$, such an optimal stopping does not necessarily exist. However, under some regularity conditions on the discounted reward process ${Z=\{Z_{t}\}_{0\leq t\leq T}}$, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, the existence and the form of an optimal stopping time are well-known from the theory of optimal stopping for random processes in continuous time, see for instance Karatzas and Shreve (1998) (Appendix D, Theorem D.12), Peskir and Shiryaev (2006) (Theorem 2.2), Lamberton (1998) (Section 10.2.1), Lamberton (2009) (Theorem 2.3.5 and Section 2.3.4), El Karoui (1981), Chapter 2. With the exception of El Karoui (1981), these results rely on the (almost sure) continuity of the reward process, which does not hold for the one involved in variable annuity pricing since its trajectories are discontinuous at time $T$ with positive probability. In fact the trajectories of $Z$ are upper semi-continuous.

### v3

@@ Abstract.
Keywords: Variable annuities, Optimal stopping, Surrender option, Discontinuous reward, Free-boundary problems
JEL Classification: C63, G12, G13, G22
Corresponding author. Email: Anne.MacKay@USherbrooke.ca
@@ 1. Introduction
Variable annuities (VA) are structured products sold by insurance companies that are mainly used for retirement planning. An initial premium is deposited in an investment account (or fund) whose return is linked to that of one or more risky assets. They are similar to mutual funds, but they also offer financial guarantees at the end of a pre-determined accumulation period. The embedded guarantees are akin to long-dated options, but they are funded by a periodic fee, generally set as a percentage of the investment account value, rather than being paid upfront. In this paper, we focus on a guaranteed minimum maturity benefit (GMMB), which provides the policyholder a minimum guaranteed amount at maturity of the contract. Another common feature of variable annuities contracts is that policyholders generally have the right to surrender, or lapse, their contract prior to maturity. When policyholders choose to do so, they receive the value accumulated in the investment account, reduced by a penalty for early surrender (or surrender charge). The uncertainty faced by VA providers with respect to early termination is known as surrender risk. Early surrenders can entail negative consequences such as liquidity issues. Thus, incorporating adequate surrender assumptions in the pricing of VAs is essential for risk management purposes; see Niittuinperä (2022).
@@
Different early surrender modelling approaches have been proposed in the literature (see Bauer et al. (2017) and Feng et al. (2022) for a review), from the use of utility functions, Gao and Ulm (2012), to modern statistical techniques, Zhu and Welsch (2015). Early surrenders are also often expressed as a decision taken by policyholders on a strictly rational basis, meaning that the contract is terminated as soon as it is optimal to do so from a financial perspective (see Bernard et al. (2014a), Jeon and Kwak (2018), Kang and Ziveyi (2018), MacKay et al. (2023), and Milevsky and Salisbury (2001), among others). Bauer et al. (2017) discusses the impact of other factors on policyholder behavior, such as taxes and expenses. This article gave rise to another stream of literature in which market frictions such as taxation rules are considered (Alonso-García et al. (2022), Bauer and Moenig (2023), and Moenig and Bauer (2016)), which helps to explain the discrepancies between market and model fee rates. Other recent studies include lapse and reentry strategies in their analysis (Bernard and Moenig (2019) and Moenig and Zhu (2018)), whereas Moenig and Zhu (2021) consider a third-party investor to whom the policyholder can sell her contract.
@@
The goal of this paper is to study the properties of the value function of the VA contract under the assumption that the policyholder maximizes its risk-neutral value, when the fee and the surrender penalty (or charge) are both time and state-dependent (that is, depending on the value of the VA account, see Bernard et al. (2014a)). Our setup includes, among others, the constant fee case, the state-dependent fees of Bernard et al. (2014a), Delong (2014), and MacKay et al. (2017), and the time-dependent fees of Bernard and Moenig (2019) and Kirkby and Aguilar (2023). Under this assumption, valuing a variable annuity contract is equivalent to solving an optimal stopping problem similar to pricing an American option. However, the financial guarantee being applied only at maturity in a VA contract creates a discontinuity in the reward function at maturity, which sets it apart from the continuous reward function of the American option and complicates the optimal stopping problem involved in the pricing of VAs. The impact of the surrender charge and the fee structure on the optimal surrender strategy has been studied in the Black-Scholes framework by Bernard et al. (2014a), Bernard and MacKay (2015), Bernard and Moenig (2019), MacKay (2014), MacKay et al. (2017), and Moenig and Zhu (2018), and in a more general setting by Kang and Ziveyi (2018) and MacKay et al. (2023). However, these articles approach the problem from a numerical perspective. Recently, and independently from our work, De Angelis et al. (2024) study the optimal surrender problem in VAs from a theoretical perspective. While their setting includes mortality, which ours does not, we consider a more general structure for the fee and surrender charge functions. The analysis presented in this paper is therefore significantly different from theirs, but a parallel can be drawn between some of their results and ours.
@@
Recently, Luo and Xing (2021) studied variable annuity contracts under regime-switching volatility models from an optimal stopping perspective. However, they consider a different surrender benefit, which results in a continuous reward function similar to that of a put option. Chiarolla et al. (2022) perform an analysis similar to ours, but for participating policies with minimum rate guarantee and surrender option. Participating policies are akin to variable annuities in that the premium paid by the policyholder tracks a financial portfolio, subject to a minimum rate guarantee. However, the problem they consider is fundamentally different from ours in multiple ways. i) In variable annuity contracts, guarantees are funded via periodic fees set as a percentage of the account value. This creates a discrepancy between the fee amount and the value of the financial guarantee, which becomes an incentive for the policyholder to surrender the policy early (Milevsky and Salisbury (2001)). In participating policies, there is no such ongoing fee; upon termination or at maturity, the policyholder receives the reserve and a given percentage of the surplus, defined as a percentage of the tracking portfolio over the policy reserve (the intrinsic value). ii) In variable annuities, the surrender value differs from that of the maturity payout, which creates a time discontinuity in the reward function at maturity; whereas in participating policies, the intrinsic value is paid at maturity or upon early termination, so that the reward function is independent of time and continuous. iii) In participating policies, the contract is terminated if the underlying portfolio falls below the reserve. There is no such feature in a variable annuity, that is, early termination is at the sole discretion of the policyholder. The resulting optimal stopping problem we study in this paper is thus very different from the one considered by Chiarolla et al. (2022), even if they are motivated by similar insurance products.
@@
The shape and the (non-)emptiness of the exercise region of an American option with a time-homogeneous reward function has first been studied rigorously by Villeneuve (1999). Other authors, such as Jönsson et al. (2006), Kotlow (1973), and Villeneuve (2007), also studied the conditions under which the exercise region has a particular shape. These results are however not directly applicable to the problem we study in this paper, because of the time-dependence and the discontinuity of our reward function. Under the assumption of a general fee and surrender charge structure, we are nonetheless able to identify a condition, expressed as a partial differential inequality, under which the optimal stopping time is always at maturity of the contract, thus mitigating surrender risk. This condition provides a powerful tool to understand the interplay between the fees and surrender charges, and its effect on the optimal stopping strategy. It can also be used to further study and characterize the shape of the optimal exercise, or surrender, region. In particular, we show that the surrender region can be a disconnected set and that the optimal surrender boundary can be discontinuous, thus illustrating a major difference between the pricing of VA contracts and standard American options. Our results also extend the work of Milevsky and Salisbury (2001), who consider an infinite horizon, to a more general finite horizon problem.
@@
The main contributions of this paper are summarized below.
@@
  • We present an alternative representation for the value function of the VA contract in terms of a continuous reward function. As stated above, various results in the literature, in particular those regarding continuity of the value function, rely on the continuity of the reward function. In our paper, this alternative representation is used in Section 4 to perform a rigorous theoretical study of the value function of the VA contract and to show that it solves a free boundary value problem. These results justify the use of different numerical methods already applied to VA pricing in the actuarial literature.
  • We provide two further representations for the value function. For the first one, we decompose the value of the contract into the value of the maturity benefit and an integral term akin to the early exercise premium from the American option literature. The second representation is new to the actuarial and American option pricing literature and expresses the contract value as the sum of the surrender benefit and an integral term coined the continuation premium. These two representations are used in Section 4.3 to better understand the value function of a VA contract and in the proofs of Section 4.4, which presents characterizations of the (non-) emptiness of the surrender region. The two representations can also be used to develop numerical pricing algorithms.
  • We obtain conditions under which the original optimal stopping problem and the one with a continuous reward function are equivalent.
@@ 2.1. Market Model
On a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, let $W=\{W_{t}\}_{t\geq 0}$ be a standard Brownian motion whose augmented filtration is denoted by $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$. $\mathbb{Q}$ is the probability measure used to price the assets presented below.
@@
where $r$, $\sigma$ are deterministic constants with $\sigma>0$. It is easy to verify that (1) has a unique strong solution given by $B_{t}=B_{0}e^{rt}$ and
@@
Readers will recognize the so-called Black-Scholes model, presented here directly in terms of its unique risk-neutral measure.
@@ 2.2. Variable Annuity Contract
We consider a simplified variable annuity contract offering a guaranteed minimum accumulation benefit at maturity $T\in\mathbb{R}_{+}$, with $\mathbb{R}_{+}$ denoting the strictly positive real numbers, that is +=$\mathbb{R}_{+}=\{x\in\mathbb{R}|x>0\}$. At inception of the contract, the policyholder deposits an initial premium in an investment account (or fund) tracking the financial market. Here, we assume that this account tracks the risky asset with price process $S$. We denote by $F=\{F_{t}\}_{0\leq t\leq T}$ the value process of the account.
We remark that variable annuity policyholders usually benefit from a return-of-premium guarantee if they die before maturity. Under the assumption that mortality risk is completely diversifiable, it is straight-forward to add the death benefit in the analysis of the contract, see for example MacKay et al. (2017). See also De Angelis et al. (2024) for a more in-depth analysis. In this paper, we ignore mortality risk.
The guarantees embedded in VA contracts are funded via a fee levied continuously from the investment account at a rate $c_{t}$ defined by
@@
where $C:[0,T]\times\mathbb{R}_{+}\rightarrow[0,1]$, so that (4) has a unique strong solution. This fee structure is general enough to include state-dependent (Bernard et al. (2014b), Delong (2014), MacKay et al. (2017)) and time-dependent fees (Bernard and Moenig (2019), Kirkby and Aguilar (2023)). The account value is then defined by
@@
At maturity, the policyholder receives the maximum between a pre-determined amount $G\in\mathbb{R}_{+}$ and the value of the investment account. Given $F_{t}=x$, the time-$t$ risk-neutral value of the maturity benefit is given by
@@
Should she decide to surrender her contract prior to maturity, the policyholder receives the amount accumulated in the account subject to a penalty, expressed as a percentage of the account, which can depend on time and the value of the account. If no surrenders occur, the maturity benefit is paid at $T$.
@@
where $g:[0,\,T]\times\mathbb{R}_{+}\rightarrow(0,1]$ is in $C^{1,2}$ with bounded derivatives on $[0,T]\times\mathbb{R}_{+}$, is non-decreasing in $t$ and satisfies $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$. This function represents the amount received by the policyholder upon surrender at $t<T$ or at maturity $T$, given that the account has value $x$.
In the literature, $1-g$ is often called the surrender charge. It is non-increasing in time; a later surrender will yield a higher proportion of the account value. In this paper, we use the term surrender charge function to refer to either $g$ or $1-g$.
Imposing $g(T,\cdot)=1$ is common in the actuarial literature. It allows the function $g$ to be defined on the closed interval $[0,T]$ and represents the fact that at maturity, the policyholder is entitled to the full amount accumulated in the account. In practice, it is also common to see the surrender charge vanish before maturity. Examples of surrender charges are listed in Palmer (2006).
@@
For $x<G$, the function $t\mapsto\varphi(t,x)$ is discontinuous at $T$ since
@@
Under the assumption that the policyholder maximizes the risk-neutral value of her contract, its value at time $t$ with account value $x$ is given by
(7) $v(t,\,x)=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}[e^{-r(\tau-t)}\varphi(\tau,F_{\tau}^{t,\,x})]=\sup_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}[e^{-r(\tau-t)}\varphi(\tau,F_{\tau}^{t,\,x})],$
for $(t,x)\in[0,T]\times\mathbb{R}_{+}$, where $\mathcal{T}_{t,\,T}$ is the set of all stopping times taking value in the interval $[t,\,T]$. Moreover, throughout the paper, it must be understood that for $s>t$,
$v(s,F^{t,x}_{s})=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,\,T}}\mathbb{E}[e^{-r(\tau-s)}\varphi(\tau,F_{\tau}^{t,\,x})],$
since in that case, the essential supremum does not simplify.
The difference between the value of the full variable annuity contract and the present value of the maturity benefit is the value of the surrender right, which we denote by $e:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}$ and define as
@@ 3. Optimal Stopping Time
In this case, we also say that $\tau^{\star}$ is optimal for (8).
For a general reward function $\varphi$, such an optimal stopping time does not necessarily exist. However, under some regularity conditions on the discounted reward process ${Z=\{Z_{t}\}_{0\leq t\leq T}}$, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, the existence and the form of an optimal stopping time are well known from the theory of optimal stopping for random processes in continuous time, see, for instance, Karatzas and Shreve (1998) (Appendix D, Theorem D.12), Peskir and Shiryaev (2006) (Theorem 2.2), Lamberton (1998) (Section 10.2.1), Lamberton (2009) (Theorem 2.3.5 and Section 2.3.4), El Karoui (1981), Chapter 2. However, with the exception of El Karoui (1981), these results rely on the (almost sure) continuity of the reward process, which does not hold for the one involved in variable annuity pricing since its trajectories are discontinuous at time $T$ with positive probability. In fact the trajectories of $Z$ are upper semi-continuous. Thus, additional work is required to show the existence of an optimal stopping time in our setting.